What is the hydroxide ion concentration in an aqueous solution with ph =6.19 ?
1
Expert's answer
2011-05-05T12:55:03-0400
pH + pOH = 14
pOH = 14 - 6.19 = 7.81
Thus the hydroxide ion concentration is 10-7.81 = 1.55 x 10-8 M